AB INITIO POTENTIAL ENERGY SURFACE AND EIGENVALUES OF A^{2}Δ STATE OF ArCH^{a}

ALEXEY L. KALEDIN, \underline{ANATOLY V. KOMISSAROV} and MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Rotationally resolved spectra of A-X transition of ArCH/D van der Waals complex have been reported previously and 7 bands were tentatively assigned. In order to make more definite assignments high level ab initio calculations have been performed.

2-D interaction potential for CH(A)-Ar (with CH bond fixed at r_{e}=1.102Å for A^{2}Δ) was calculated with MRSDCI method including Davidson and counterpoise correction. The basis set was Dunning’s correlation-consistent vqz with diffuse s and p orbitals.

The potential surface exhibits two minima, linear Ar...H-C and Ar...C-H whose respective well depths are 85 cm^{−1} and 60 cm^{−1}. The potential anisotropy is dominated by dipole and quadrupole terms on CH. Bound ro-vibrational eigentates were calculated for various J values and compared with experimental results.

aWork supported by National Science Foundation